# Weak–strong uniqueness for bi-fluid compressible system with algebraic closure

### v1

@@ 3.2 Derivation of relative entropy inequality
$\int_{\Omega}\alpha\left({H_{+}}^{\prime}({\widetilde{\varrho}_{+}}){\widetilde{\varrho}_{+}}-{H_{+}}({\widetilde{\varrho}_{+}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{\Omega}\alpha{p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}(\alpha{p_{+}}({\widetilde{\varrho}_{+}}))\ {\rm d}{x}\,{\rm d}t.$
@@
$\int_{\Omega}(1-\alpha)\left({H_{-}}^{\prime}({\widetilde{\varrho}_{-}}){\widetilde{\varrho}_{-}}-{H_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}((1-\alpha){p_{-}}({\widetilde{\varrho}_{-}}))\ {\rm d}{x}\,{\rm d}t.$
Step 6. Summing the above expressions we obtain all together:
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)\ {\rm d}{x}\,{\rm d}t-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}p\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\varrho_{+}}(\partial_{t}+{\bf u}\cdot\nabla){H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}(\partial_{t}+{\bf u}\cdot\nabla){H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\,{\rm d}t.\end{split}$ (3.17)
@@
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}-\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)+{\bf S}(\nabla{\bf v}):\nabla({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}p\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\varrho_{+}}(\partial_{t}+{\bf u}\cdot\nabla){H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}(\partial_{t}+{\bf u}\cdot\nabla){H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\,{\rm d}t.\end{split}$ (3.18)
@@
$\quad=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(\nabla\alpha{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\widetilde{\varrho}_{+}}\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t$
@@
and similarly
$-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\,{\rm d}t$
$\quad=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}\alpha{p_{+}}({\widetilde{\varrho}_{+}})+\partial_{t}(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t$
$\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\widetilde{\varrho}_{+}}\partial_{t}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
@@
Substituting for these terms in (LABEL:sumstep1b), we obtain
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}-\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big((R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)+{\bf S}(\nabla{\bf v}):\nabla({\bf v}-{\bf u})\Big)\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}})\partial_{t}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}})\partial_{t}{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})-p\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}{\bf v}-{\varrho_{+}}{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}{\bf v}-{\varrho_{-}}{\bf u})\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\nabla\alpha{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\Big)\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\partial_{t}\alpha{p_{+}}({\widetilde{\varrho}_{+}})+\partial_{t}(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t\\ &=:I+II+III+\ldots+VII+VIII.\end{split}$ (3.19)
This is an analogous form of the relative entropy inequality as the one derived in Theorem 5.2 from [15]. Let us stress that up until now we did not use any properties of $(\beta,{\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}},{\bf v})$ other than the sufficient regularity in order to use them as the test functions.
@@ 4 Relative entropy inequality for the strong solution
We will now restrict our class of test functions to $(\beta,{\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}},{\bf v})$ given in Theorem 2.3. We reformulate and reduce each of the terms from (LABEL:sumstep1c). In particular,
@@
This means we have
@@
Finally, for the terms $VII$ and $VIII$ we keep them at the present form.
@@
$\begin{split}&{\cal R}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left[(R+Q-{\widetilde{R}}-{\widetilde{Q}})\partial_{t}{\bf v}+\left((R+Q){\bf u}-({\widetilde{R}}+{\widetilde{Q}}){\bf v}\right)\cdot\nabla{\bf v}\right]\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha\Big({p_{+}}({\widetilde{\varrho}_{+}})-{p_{+}}^{\prime}({\widetilde{\varrho}_{+}})({\widetilde{\varrho}_{+}}-{\varrho_{+}})-{p_{+}}({\varrho_{+}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)\Big({p_{-}}({\widetilde{\varrho}_{-}})-{p_{-}}^{\prime}({\widetilde{\varrho}_{-}})({\widetilde{\varrho}_{-}}-{\varrho_{-}})-{p_{-}}({\varrho_{-}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})(\alpha{\varrho_{+}}-\beta{\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\Big((1-\alpha){\varrho_{-}}-(1-\beta){\widetilde{\varrho}_{-}}\Big)\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(\partial_{t}\alpha+{\bf v}\cdot\nabla\alpha\right){p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\partial_{t}(1-\alpha)+{\bf v}\cdot\nabla(1-\alpha)\Big){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\Big(\nabla\beta{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\beta){p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}(1-\beta)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}\beta\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &=:J_{1}+\ldots+J_{10}.\end{split}$ (4.1)

### v2

@@ 3.2 Derivation of relative entropy inequality
Note that (3.13) is not precisely the weak formulation; however, it can be easily deduced by considering the test function $\frac{1}{2}|{\bf v}|^{2}\psi_{\varepsilon}(t)$ with $\psi_{\varepsilon}\in C^{\infty}_{c}([0,\tau))$, $\psi_{\varepsilon}\to{\mathbf{1}}_{[0,\tau]}$ uniformly for $\varepsilon\to 0_{+}$. The same also applies below.
@@
$\int_{\Omega}\alpha\left({H_{+}}^{\prime}({\widetilde{\varrho}_{+}}){\widetilde{\varrho}_{+}}-{H_{+}}({\widetilde{\varrho}_{+}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{\Omega}\alpha{p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\Big|_{0}^{\tau}.$
@@
$\int_{\Omega}(1-\alpha)\left({H_{-}}^{\prime}({\widetilde{\varrho}_{-}}){\widetilde{\varrho}_{-}}-{H_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{\Omega}(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\Big|_{0}^{\tau}.$
Step 6. Summing the above expressions, we obtain all together:
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\int_{\Omega}(R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)\ {\rm d}{x}\,{\rm d}t-\int_{0}^{\tau}\!\!\int_{\Omega}p\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\int_{\Omega}\alpha{\varrho_{+}}(\partial_{t}+{\bf u}\cdot\nabla){H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}(\partial_{t}+{\bf u}\cdot\nabla){H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{\Omega}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}.\end{split}$ (3.17)
@@
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}-\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\int_{\Omega}(R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)+{\bf S}(\nabla{\bf v}):\nabla({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\int_{\Omega}p\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\int_{\Omega}\alpha{\varrho_{+}}(\partial_{t}+{\bf u}\cdot\nabla){H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}(\partial_{t}+{\bf u}\cdot\nabla){H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{\Omega}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}.\end{split}$ (3.18)
@@
$\quad=\int_{0}^{\tau}\!\!\int_{\Omega}\nabla{p_{+}}({\widetilde{\varrho}_{+}})\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\quad=\int_{0}^{\tau}\!\!\int_{\Omega}\Big(\alpha\nabla{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha)\nabla{p_{-}}({\widetilde{\varrho}_{-}})\Big)\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\quad=\int_{0}^{\tau}\!\!\int_{\Omega}\alpha{\widetilde{\varrho}_{+}}\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t$
@@
due to equality of the pressures, and similarly
$\int_{\Omega}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}$
$\quad=\int_{0}^{\tau}\!\!\int_{\Omega}\Big(\alpha\partial_{t}{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha)\partial_{t}{p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t$
$\quad=\int_{0}^{\tau}\!\!\int_{\Omega}\alpha{\widetilde{\varrho}_{+}}\partial_{t}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
@@
Substituting for these terms into (LABEL:sumstep1b), we obtain
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}-\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\int_{\Omega}\Big((R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)+{\bf S}(\nabla{\bf v}):\nabla({\bf v}-{\bf u})\Big)\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}})\partial_{t}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}})\partial_{t}{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}\Big(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})-p\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}{\bf v}-{\varrho_{+}}{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}{\bf v}-{\varrho_{-}}{\bf u})\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &=:I+II+III+\ldots+VI.\end{split}$ (3.19)
[…]
@@ 4 Relative entropy inequality for the strong solution
We will now restrict our class of test functions to $(\beta,{\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}},{\bf v})$ given in Theorem 2.4, in particular satisfying (2.13). We reformulate and reduce each of the terms from (LABEL:sumstep1c). Then
@@
This means that we have
@@
$\begin{split}&{\cal R}=\int_{0}^{\tau}\!\!\int_{\Omega}\left[(R+Q-{\widetilde{R}}-{\widetilde{Q}})\partial_{t}{\bf v}+\left((R+Q){\bf u}-({\widetilde{R}}+{\widetilde{Q}}){\bf v}\right)\cdot\nabla{\bf v}\right]\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}\alpha\Big({p_{+}}({\widetilde{\varrho}_{+}})-{p_{+}}^{\prime}({\widetilde{\varrho}_{+}})({\widetilde{\varrho}_{+}}-{\varrho_{+}})-{p_{+}}({\varrho_{+}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}(1-\alpha)\Big({p_{-}}({\widetilde{\varrho}_{-}})-{p_{-}}^{\prime}({\widetilde{\varrho}_{-}})({\widetilde{\varrho}_{-}}-{\varrho_{-}})-{p_{-}}({\varrho_{-}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})(\alpha{\varrho_{+}}-\beta{\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\Big((1-\alpha){\varrho_{-}}-(1-\beta){\widetilde{\varrho}_{-}}\Big)\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\Big(\nabla\beta{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\beta){p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}(1-\beta)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}\beta\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &=:J_{1}+\ldots+J_{8}.\end{split}$ (4.1)